OpenStudy (anonymous):

Solve log2x-log(x-2)=0. a.) 8 b.) -1 c.) no solution d.) 2.5

OpenStudy (anonymous):

start with \[\log(\frac{2x}{x-2})=0\]

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

then rewrite in equivalent exponential form a s \[\frac{2x}{x-2}=10^0\] or just \[\frac{2x}{x-2}=1\] and solve for \(x\)

OpenStudy (anonymous):

be careful here let me know what you get

OpenStudy (anonymous):

there is a trap set here can you solve the equation i wrote above?

OpenStudy (anonymous):

im trying to.

OpenStudy (anonymous):

i cant do it on the calculator.

OpenStudy (anonymous):

\[\frac{2x}{x-1}=1\]is not a calculator exercise start with \[2x=x-2\] and solve for \(x\)

OpenStudy (anonymous):

is it 2?

OpenStudy (anonymous):

or just 2x=x

OpenStudy (anonymous):

no it is \(-2\) which means there is no solution because you cannot take the log of -2

OpenStudy (anonymous):

oh. well thank you!

OpenStudy (anonymous):

yw